OpenStudy (anonymous):

hi :) 1(a^8) + 4(a^4)b +4(b^2)-1(c^4) factor this to products of two binomial

OpenStudy (mathstudent55):

\( 1(a^8) + 4(a^4)b +4(b^2)-1(c^4) \) \( = a^8 + 4a^4b +4b^2-c^4 \) \(= a^8 + (4b)a^4 + (2b + c^2)(2b - c^2) \) The expression above is of the form \(x^2 + bx + c\) where \(a = 1\) \(b = 4b\), and \(c = (2b + c^2)(2b - c^2)\) To factor it, we need to factors of c that add to b. We have two factors of c already: \(2b + c^2\) and \(2b - c^2\) We need the factors to add to the middle coefficient which is 4b. \(2b + c^2 + 2b - c^2= 4b\) It works, so we can factor: \(= a^8 + (4b)a^4 + (2b + c^2)(2b - c^2) \) \(= (a^4 + 2b + c^2)(a^4 + 2b - c^2)\) This was factored into two trinomials, not two binomials.

OpenStudy (mathstudent55):

\( 1(a^8) + 4(a^4)b +4(b^2)-1(c^4) \) \( = a^8 + 4a^4b +4b^2-c^4 \) \(= a^8 + (4b)a^4 + (2b + c^2)(2b - c^2) \) The expression above is of the form \(x^2 + bx + c\) where \(a = 1\) \(b = 4b\), and \(c = (2b + c^2)(2b - c^2)\) To factor it, we need to factors of c that add to b. We have two factors of c already: \(2b + c^2\) and \(2b - c^2\) We need the factors to add to the middle coefficient which is 4b. \(2b + c^2 + 2b - c^2= 4b\) It works, so we can factor: \(= a^8 + (4b)a^4 + (2b + c^2)(2b - c^2) \) \(= (a^4 + 2b + c^2)(a^4 + 2b - c^2)\) This was factored into two trinomials, not two binomials.